# The anonymization problem in social networks

### v1

@@ 3.2.1 Alteration operations
Vertex Refinement Query [37, 21] vrq $\{degree(u):u\in V,\ dist(u,v)\leq d\}$ Degree iterative $V_{N_{d}(v)}\cup V_{N_{d}(w)}$
3.2.2 Measures for k-anonymity
Various measures for $k$-anonymity have been introduced in the literature. As described in [25], the choice of measure essentially determines the attacker scenario one protects against and has an effect on the uniqueness of the network. In this paper, we use the measures listed in Table 1. We choose one measure for each category in Section 2, hence representing the most common measures used in the literature. While for completeness sake measures are parameterized by their reach, i.e., distance $d$, we focus on $d=1$ in this paper.
3.2.3 Affected nodes
During the anonymization process we can assign a state to each node equal to the value computed by the anonymity measure. After an edge is altered, the state of a number of nodes changes. Which nodes are affected depends on the reach of the chosen measure, i.e., up to where structural information is taken into account to compute the anonymity of a node [25]. We let $A_{M}(\{v,w\})$ in the rightmost column of Table 1 denote the set of nodes affected by altering edge $\{v,w\}$ using anonymity measure $M$.
@@ 4 The anonymization problem
Definition 1 (Anonymization problem.)
Given a network $G$, a set of allowed operations, a measure for anonymity $M$, and value for $k$:
@@
    Full anonymization: Make all nodes $k$-anonymous with as few alterations as possible.
@@
    Partial anonymization: Ensure that $x\%$ of the nodes is $k$-anonymous with as few alterations as possible.
@@
    Budgeted anonymization: Alter the network with at most $b$ operations to make as many nodes as possible $k$-anonymous.
Most commonly studied is the first variant which aims to make the entire network $k$-anonymous. However, since some nodes may require more work to anonymize, we introduce the partial variant aiming to anonymize a certain fraction of the nodes. This generalizes the subset variant proposed in  [42] aiming to anonymize a subset of nodes. Additionally, the budgeted variant of the problem allows up to $b$ edge alterations. This enables the user to balance the privacy/utility trade-off, as generally with fewer alterations, more network utility is preserved.
In the remainder of this paper we aim to better understand the anonymization problem and its three variants under different settings. Therefore, in our experiments we compare results using various operations or measures for anonymity, and the performance of heuristic based anonymization algorithms on the three variants. For the budgeted variant, we use a budget of 5% as work on robustness of network suggests that randomly altering 5% of the edges preserves high utility overall [35]. We focus on $k=2$ and measures with $d=1$, as previous work [7] showed that anonymity does not change drastically for $2<k\leq 5$. Moreover, for $d\geq 2$ many more if not all nodes quickly become unique, whereas this does not per se represent a realistic attacker scenario.
@@ 5 Approach
In this section we first present the general framework for anonymization. Then, we introduce the seven (one baseline, six new) measure-agnostic algorithms used in the remainder of this paper.
@@ 5.1 Anonymization
To anonymize a network, we propose Algorithm 1. This algorithm can be used for all variants of the anonymization problem introduced in Section 4, based on the target anonymity and budget given as input. For the full and partial variant, the budget should equal $|E|$ and the target anonymity parameter anon_stop should be set to $|V|$ and $x*|V|$, respectively. For the budgeted variant, one should give the budget and a target anonymity of $|V|$. The other inputs required are the original graph, the partition of nodes computed, the chosen measure (Table 1), the anonymization algorithm to select edges and the recompute gap, allowing for lazy evaluation (see Secton 5.2).
1 Input: graph $G$, partition $P_{M}$, measure $M$, value of $k$, anonymization algorithm $Anonalg$, budget $b$, target anonymity $anon\_stop$, recompute gap $recompute\_gap$
[…]
3while $b>0$ and $anonymity(G^{\prime},k)>anon\_stop$ do
4      $nr\_alter\leftarrow min(recompute\_gap,b)$
5       $altered\_edges\leftarrow Anonalg(G^{\prime},P_{M},M,k,nr\_alter)$
6       $G^{\prime}\leftarrow update\_graph(G^{\prime},altered\_edges)$
7       $P_{M}\leftarrow update\_P(G^{\prime},P_{M},M,altered\_edges)$
8       $b\leftarrow b-|altered\_edges|$
9       if $anonymity(G^{\prime},k)>anonymity(G_{best},k)$ then
10             $G_{best}\leftarrow G^{\prime}$
11       end if
12
13 end while
Return $G_{best}$
@@ 10             $G_{best}\leftarrow G^{\prime}$
The algorithm works as follows. During each iteration of the while loop in lines  1 to 1, a set of edges is altered. In line 1 this set is chosen by the selected anonymization algorithm (see Section 5.3). The network is altered in line 1 and the affected part of the partition is updated accordingly in line 1. The latter is done in an efficient manner, as described in Algorithm 2. The remainder of the lines are bookkeeping operations to ensure the right number of edges is deleted and the best result is stored. Here the anonymity() function returns how many nodes are $k$-anonymous. When the while loop terminates, either the budget is depleted or the target anonymity is reached, and the most anonymous network is returned in line 1.
5.2 Lazy evaluation
Altering the edges changes the network structure, which in turn influences the partition of nodes, and hence their anonymity. This can affect the edges chosen by anonymization algorithms if they account for the structure surrounding an edge or the uniqueness of nodes. By design, we allow for for a trade-off between accuracy and computational cost by updating the graph and recomputing the partition after recompute gap alterations, rather than after each graph alteration. A value of 1 implies the graph and partition is constantly updated, i.e., no recompute gap, while for $b/100$ the values are updated 100 times. Results in Appendix A of the Supplementary materials show that using a recompute gap of 1 results in similar anonymization performance compared to when a larger recompute gap is used, yet resulting in substantially higher runtimes.
1 Input: graph $G$, partition $P_{M}$, Measure $M$, altered edges $altered\_edges$
[…]
3 for $e\in altered\_edges$ do
4       $update\_nodes\leftarrow update\_nodes\cup A_{M}(e)$
5 end for
6
[…]
Return $P_{M}$
Algorithm 2 update_p
Algorithm 2 describes how the equivalence partition is updated given a set of altered edges. First, the for loop in lines 2 to 2 finds the set of affected nodes, i.e., the set of nodes for which the state changes (cf. the $A_{M}()$ function in the rightmost column of Table 1). For these nodes the equivalence class is updated in line 2. Lastly, the updated equivalence partition is returned.
@@ 5.3 Anonymization algorithms
In this section we describe the seven heuristic anonymization algorithms that each work independently from the anonymity measure and differ in terms of what type of information they take into account. In each heuristic, edges are selected by assigning a probability to each edge and based on these, selecting the desired number of edges nr_delete. As described in Algorithm 1, the goal is to return a set of edges to be altered, or more precisely in our case; deleted, as this was found to be the most promising operation in Section 6.2. This is also to be expected theoretically, as deleting edges overall results in neighborhoods with fewer nodes and edges. As the neighborhoods are smaller, there are fewer possible neighborhoods with the same number of nodes which in turn increases the probability of the node being unique. The deletion of an edge has a larger effect on the nodes it connects as this both removes the neighbor and all triangles formed with this node from the neighborhood. For other affected nodes one edge is deleted from the neighborhood.
The introduced algorithms can be categorized as follows: 1) edge sampling (baseline), 2) structure based, and 3) uniqueness based.
@@ 5.3.2 Structure based
This type of heuristic targets edges based on a node’s surrounding network structure. degmin does so by assigning higher probabilities if both nodes connected by the edge have a high degree. This has a high chance to affect many nodes.
$P_{\text{{degmin}}}(\{v,w\})=\frac{min(degree(v),degree(w))}{\sum_{\{v^{\prime},w^{\prime}\}\in E}min(degree(v^{\prime}),degree(w^{\prime}))}$
It can also be beneficial to target low degree nodes, as real-world networks tend to have many low degree nodes with a similar degree due to the powerlaw degree distribution. These nodes overall require fewer changes to anonymize. The degdiff algorithm aims to target edges connecting a high and low degree node.
$P_{\text{{degdiff}}}(\{v,w\})=\frac{|degree(v)-degree(w)|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}|degree(v^{\prime})-degree(w^{\prime})|}$
Lastly, aff computes the exact number of nodes affected by deleting an edge, $A_{M}(\{v,w\})$, as defined in the rightmost column of Table 1. For measures with a reach larger than the degree, this is more computationally expensive to determine ($\mathcal{O}(|V|^{2})$) than the degree of nodes which is in $\mathcal{O}(1)$.
$P_{\text{{aff}}}(\{v,w\})=\frac{|A_{M}(\{v,w\})|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}|A_{M}(\{v^{\prime},w^{\prime}\})|}$
@@ 5.3.3 Uniqueness based
Since only unique nodes need to be altered in order to increase anonymity, this category of algorithms aims to have a larger effect on unique nodes.
@@
The first algorithm, unique, filters the set of edges to only contain unique edges, i.e., edges connected to at least one unique node. We refer to the set of unique edges as $E_{u}$ and the set of unique nodes as $V_{u}$. After filtering, the desired number of edges is chosen randomly. If the set of unique edges is not large enough, the remainder of the edges is randomly selected.
$P_{\text{{unique}}}(\{v,w\})=\begin{cases}1/|E_{u}|&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|>nr\_delete$}\\ 1&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|<=nr\_delete$}\\ 1/|E-E_{u}|&\text{if $\{v,w\}\notin E_{u}$, $|E_{u}|<nr\_delete$}\\ 0&\text{otherwise}\end{cases}$
@@
The second algorithm, aff-u, takes into account how many unique nodes are affected by altering the given edge. As a result, edges affecting more unique nodes are more likely to be chosen. The factor of $\frac{1}{|E|}$ is added to avoid probabilities equal to 0.
$P_{\text{{aff-u}}}(\{v,w\})=\frac{|A_{{M}}(\{v,w\})\cap V_{u}|+1/|E|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}|A_{M}(\{v^{\prime},w^{\prime}\})\cap V_{u}|+1/|E|}$
The third algorithm, u-aff-u, combines the previous two algorithms by first filtering for unique edges (unique), and determining the probability for each edge based on how many nodes are affected (aff-u). The filtering step allows the algorithm to target edges that are expected to both have a big effect and are unique.
$P_{\text{{u-aff-u}}}(\{v,w\})=\begin{cases}P_{\text{{aff-u}}}(\{v,w\})&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|>nr\_delete$}\\ 1&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|<=nr\_delete$}\\ 1/|E-E_{u}|&\text{if $\{v,w\}\notin E_{u}$, $|E_{u}|<nr\_delete$}\\ 0&\text{otherwise}\end{cases}$
@@
Table 2: Real-world networks used for experiments and their properties: the number of nodes, edges, average degree, clustering coefficient, assortativity, diameter and average distance. Results: the initial uniqueness using count, the best (highest) values obtained for each variant of the anonymization problem using the es and heuristic algorithms; and the network category.
Frac. edges preserved
Frac. unique nodes
anonymized
1. Full 2. Partial 3. Budgeted
Network $|V|$ $|E|$
@@
Unique-
ness
ES [6] Heur. Ratio. ES [6] Heur. Ratio ES [6] Heur. Ratio Category
Radoslaw emails [45] 167 3,250 38.92 0.686 -0.295 5 1.97 0.766 0.011 0.106 9.26 0.017 0.120 6.91 0.028 0.033 1.17 Communication
Primary school [46] 236 8,317 49.99 0.502 0.173 3 1.86 0.975 0.006 0.044 7.29 0.022 0.100 4.54 0.002 0.009 5.31 Human contact
Moreno innov. [45] 241 923 7.66 0.313 -0.056 5 2.47 0.245 0.109 0.676 6.21 0.191 0.717 3.76 0.214 0.437 2.05 Communication
Gene fusion [45] 291 279 1.92 0.003 -0.355 9 3.90 0.024 0.223 0.898 4.03 0.419 0.935 2.23 0.171 0.829 4.84 Biological
Copnet calls [47] 536 621 2.32 0.255 0.172 22 7.37 0.024 0.204 0.905 4.44 0.602 0.938 1.56 0.062 0.985 16.01 Communication
Copnet sms [47] 568 697 2.45 0.223 0.192 20 7.32 0.026 0.225 0.947 4.20 0.702 0.972 1.38 0.267 1.000 3.75 Communication
Copnet FB [47] 800 6,418 16.05 0.323 0.182 7 2.98 0.488 0.015 0.290 19.63 0.128 0.501 3.90 0.185 0.28 1.51 Online social
FB Reed98 [48] 962 18,812 39.11 0.330 0.023 6 2.46 0.778 0.001 0.065 101.90 0.057 0.250 4.42 0.072 0.088 1.21 Online social
Arenas email [45] 1,133 5,451 9.62 0.254 0.078 8 3.61 0.230 0.035 0.461 13.13 0.219 0.637 2.90 0.107 0.299 2.79 Communication
Euroroads [45] 1,174 1,417 2.41 0.020 0.127 62 18.37 0.003 0.522 0.98 1.88 1.000 1.000 1.00 0.467 1.000 2.14 Roads
Air traffic control [45] 1,226 2,408 3.93 0.073 -0.015 17 5.93 0.042 0.079 0.831 10.51 0.454 0.914 2.01 0.200 0.859 4.29 Air traffic
Network science [45] 1,461 2,742 3.75 0.878 0.462 17 5.82 0.039 0.065 0.781 12.10 0.303 0.894 2.95 0.070 0.751 10.70 Co-autorship
FB Simmons81 [48] 1,518 32,988 43.46 0.325 -0.062 7 2.57 0.785 0.003 0.088 32.99 0.065 0.274 4.25 0.059 0.077 1.30 Online social
DNC emails [45] 1,866 4,384 4.70 0.587 -0.307 8 3.37 0.092 0.007 0.237 32.07 0.123 0.396 3.22 0.108 0.158 1.46 Online social
Moreno health [45] 2,539 10,455 8.24 0.151 0.251 10 4.56 0.054 0.097 0.717 7.41 0.509 0.942 1.85 0.285 0.960 3.36 Human social
US power grid [45] 4,941 6,594 2.67 0.107 0.003 46 18.99 0.008 0.131 0.943 7.22 1.000 1.000 1.00 0.297 1.000 3.36 Power grid
GRQC collab. [49] 5,241 14,484 5.53 0.687 0.659 17 6.05 0.054 0.022 0.423 19.49 0.242 0.747 3.08 0.006 0.065 10.26 Co-autorship
6 Results
In this section we discuss the results of experiments aiming to understand the anonymization problem. First, we summarize the experimental setup and network datasets used. Second, we discuss experiments comparing the effectiveness of different alteration operations on anonymization of graph models. Third, we move to real-world networks and compare the effect of using different anonymity measures. Fourth, we compare the anonymization algorithms based on their effectiveness on the three variants of the problem. Lastly, we compare the runtimes of the algorithms.

### v2

@@ 3.2.2 Measures for k-anonymity
Vertex Refinement Query [43, 22] vrq
$\{degree(u):u\in V,$
$dist(u,v)\leq d\}$
Degree iterative $V_{N_{d}(v)}\cup V_{N_{d}(w)}$
After an edge is altered, the state of a number of nodes changes. Which nodes are affected depends on the reach of the chosen measure, i.e., up to where structural information is taken into account to compute its state [26]. We let $A_{M}(\{v,w\})$ in the rightmost column of Table 1 denote the set of nodes affected by altering edge $\{v,w\}$ using anonymity measure $M$. The reach depends on the type of measure considered, and can be categorized into degree based, neighborhood based and based on the degree of nodes that are iteratively further away. Computing the set of affected nodes is computationally least expensive for the degree and iterative degree based measures. For these measures, the affected nodes are the two nodes an edge $\{v,w\}$ connects, and the union of all nodes at distance $d$ from nodes $v,w$, respectively. Determining these sets of nodes requires a computational complexity of $\mathcal{O}(1)$ and $\mathcal{O}(|E|)$, respectively. For the neighborhood type measures, a node $u$ is affected if the edge $\{v,w\}$ occurs in $N_{d}(u)$. To determine the set of affected nodes, the intersection of the sets $V_{N_{d}}(v)$ and $V_{N_{d}}(w)$ needs to be determined, resulting in a computational complexity of $\mathcal{O}(|E|^{2})$.
@@ 4 The anonymization problem
Definition 1 (Anonymization problem.).
Given a network $G$, a measure for anonymity $M$, value for $k$ and one or more allowed alteration operations:
@@
    Full anonymization: Make all nodes $k$-anonymous with as few alterations to the network as possible.
@@
    Partial anonymization: Ensure that a chosen fraction $0.0\leq\alpha\leq 1.0$ of the nodes is $k$-anonymous with as few alterations to the network as possible.
@@
    Budgeted anonymization: Alter the network with at most $B$ alterations to the network to make as many nodes as possible $k$-anonymous.
Most commonly studied is full anonymization which aims to make the entire network $k$-anonymous. However, since some nodes may require more work to anonymize, we introduce the partial variant aiming to anonymize a certain fraction of the nodes. This generalizes the subset variant proposed in [48] for which the aim is to anonymize a given subset of nodes. Additionally, the budgeted variant of the problem allows up to $B$ edge alterations. This enables the user to balance the anonymity versus utility trade-off, as generally with fewer alterations, more network utility is preserved.
In the remainder of this paper we aim to better understand the anonymization problem in its three variants under different settings. Therefore, in our experiments, we first perform two preliminary experiments showing how uniqueness changes when applying edge perturbations while using various 1) alteration operations and 2) anonymity measures. Second, we compare the performance of heuristic based anonymization algorithms on the three variants of the problem and compare how the algorithms preserve utility. For the budgeted variant, we use a budget of 5% as work on robustness of network suggests that randomly altering 5% of the edges preserves a high utility overall [41, 54]. We focus on $k=2$ and measures with $d=1$, as for $d\geq 2$ many more, if not all nodes, quickly become unique, whereas this does not per se represent a realistic attacker scenario.
@@ 5 Approach
In this section we first present the general framework for anonymization. Then, we introduce the five (one baseline, four new) measure-agnostic algorithms used in the remainder of this paper.
@@ 5.1 Anonymization
To anonymize a network, we propose the framework outlined in Algorithm 1. This general framework can be used for all variants of the anonymization problem introduced in Section 4 based on the target anonymity, that is, the desired number of nodes that is at least $k$-anonymous, and budget given as input. For the full and partial variant, the budget should equal $|E|$ and the target anonymity parameter $T$ should be set to $|V|$ and $\alpha*|V|$ respectively, where $\alpha$ equals the fraction of nodes that should be anonymous for the partial variant of the problem. For the budgeted variant, one should give the budget $B$ and a target anonymity of $|V|$. The other inputs required are the original graph $G$, the partition $P_{M}$ of nodes computed using the chosen measure (Table 1), the anonymization algorithm, and recompute gap $R$, which determines after how many edge alterations the equivalence partition should be recomputed. We discuss the use and effect of this parameter on runtime and performance in Section 5.2.
1:Input: graph $G$, partition $P_{M}$, measure $M$, value of $k$, anonymization algorithm $Anonalg$, budget $B$, target anonymity $T$, recompute gap $R$
2:$G^{\prime}\leftarrow G$, $G_{best}\leftarrow G$
3:while $B>0$ and $anonymity(G^{\prime},P_{M},k)<T$ do
4:     $B^{\prime}\leftarrow min(R,B)$
5:     $E_{A}\leftarrow Anonalg(G^{\prime},P_{M},M,k,B^{\prime})$triangleright See section 5.3
6:     $G^{\prime}\leftarrow update\_graph(G^{\prime},E_{A})$
7:     $P_{M}\leftarrow update\_P_{M}(G^{\prime},P_{M},M,E_{A})$triangleright Algorithm 2
8:     $B\leftarrow B-B^{\prime}$
9:     if $anonymity(G^{\prime},P_{M},k)>anonymity(G_{best},P_{M},k)$ then
10:         $G_{best}\leftarrow G^{\prime}$
11:     end if
12:end while
13:Return: $G_{best}$
@@
The algorithm works as follows. During each iteration of the while loop in lines  3 to 12, a set of edges is altered. In line 5 this set is chosen by the selected anonymization algorithm (see Section 5.3). The network is altered in line 6 and the affected part of the partition is updated accordingly in line 7. The latter is done in an efficient manner, described in Algorithm 2, to avoid recomputation of node states and anonymity for the entire graph, which is a computationally expensive part of the algorithm. The remainder of the lines are bookkeeping operations to ensure the right number of edges $B^{\prime}$ is deleted each iteration and the best result, $G_{best}$, is updated. The anonymity() function returns how many nodes are $k$-anonymous in the altered graph. When the while loop terminates, either budget $B$ is depleted or target anonymity $T$ is reached. Finally, line 13 returns the graph with the highest anonymity found by the algorithm.
5.2 Recompute gap
Altering the edges changes the network structure, which in turn influences the state of nodes, and hence their anonymity. This can affect the edges chosen by anonymization algorithm if this accounts for the structure surrounding an edge, or the uniqueness of nodes. As updating the equivalence classes and datastructures used by anonymization algorithms is computationally expensive, we allow for a trade-off between accuracy and computational cost by updating the graph and recomputing the partition after $R$ alterations, rather than after each graph alteration. A recompute gap value of 1 implies that the graph and partition are updated after each edge deletion, i.e., no recompute gap, while for $B/100$ the values are updated 100 times. Results in Appendix A show that using a recompute gap of 1 results in similar anonymization performance compared to when a larger recompute gap is used, yet resulting in substantially higher runtimes.
1:Input: graph $G$, partition $P_{M}$, Measure $M$, altered edges $E_{a}$
2:$V_{a}=\bigcup_{e\in E_{a}}A_{M}(e)$triangleright Set of affected nodes (Table 1)
3:for $v\in V_{a}$ do
4:     $state=M(v)$triangleright New node state
5:     $P_{M}=update(P_{M},v,state)$
6:end for
7:Return: $P_{M}$
Algorithm 2 update_pM
Algorithm 2 describes how the equivalence partition is updated given a set of altered edges. First, line 2 finds the set of affected nodes, i.e., the set of nodes for which the state changes cf. the $A_{M}()$ function in the rightmost column of Table 1. For these nodes the state is computed in lines 4, and the equivalence classes are updated accordingly in line 5. Lastly, the updated equivalence partition is returned.
@@ 5.3 Anonymization algorithms
In this section we describe the five heuristic anonymization algorithms. Each of them works independently from the considered anonymity measure and differs in terms of what type of information they take into account. For each heuristic, edges are selected by assigning a probability to each edge and based on these, selecting the desired number of edges. As described in Algorithm 1, the goal is to return a set of edges to be altered, or more precisely in our case; deleted, as this was found to be the most promising operation in Section 6.2 compared to edge addition and rewiring.
The edge deletion operation is likely the most effective as deleting edges decreases the number of nodes and edges in the neighborhood of nodes. This in turn results in smaller neighborhoods for which fewer network structures are possible and it is less likely to be unique, possibly making more nodes anonymous. The deletion of an edge has a larger effect on the states of the nodes it connects as this removes for both nodes one neighbor, and all triangles formed with this node from the neighborhood. For other affected nodes one edge is deleted from the neighborhood. In our heuristic algorithms, we account for these findings.
The five proposed heuristic algorithms can be categorized as follows: 1) edge sampling (baseline), 2) structure based, and 3) uniqueness based.
@@ 5.3.2 Structure based
These heuristics target edges based on a node’s surrounding network structure. The degree heuristic does so by assigning higher probabilities if both nodes connected by the edge have a high degree. By targeting these edges, this heuristic is likely to affect more nodes.
$P_{\text{{degree}}}(\{v,w\})=\frac{min(degree(v),degree(w))}{\sum_{\{v^{\prime},w^{\prime}\}\in E}min(degree(v^{\prime}),degree(w^{\prime}))}$ (6)
The aff heuristic computes the exact number of nodes affected by deleting an edge as defined in the rightmost column of Table 1. Computing the number of affected nodes for the count anonymity measure, which we focus on in this paper, has a computational complexity of $\mathcal{O}(|V|^{2})$, as discussed in Section 3.2.2. This makes the heuristic value more expensive to compute compared to the degree heuristic.
$P_{\text{{aff}}}(\{v,w\})=\frac{|A_{M}(\{v,w\})|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}|A_{M}(\{v^{\prime},w^{\prime}\})|}$ (7)
@@ 5.3.3 Uniqueness based
Since primarily unique nodes need to be altered in order to increase anonymity, this category of algorithms aims to have a larger effect on unique nodes.
@@
The first uniqueness based algorithm, unique, is guaranteed to change the state of unique nodes, as it only targets the set of unique edges, i.e., edges connected to at least one unique node, as denoted in Equation 4. From this set, the desired number of edges $B^{\prime}$ is chosen at random. If the set of unique edges is not large enough, the remainder of the edges is randomly selected from the set of all edges in the graph.
$P_{\text{{unique}}}(\{v,w\})=\begin{cases}1/|E_{u}|&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|>B^{\prime}$}\\ 1&\text{if $\{v,w\}\in E_{u}$, $|E_{u}|\leq B^{\prime}$}\\ 1/|E-E_{u}|&\text{if $\{v,w\}\notin E_{u}$, $|E_{u}|<B^{\prime}$}\\ 0&\text{otherwise}\end{cases}$ (8)
@@
The second uniqueness based algorithm, ua, takes into account the number of unique nodes that are affected by altering the given edge. As a result, edges affecting more unique nodes are more likely to be chosen. The factor of $\frac{1}{|E|}$ is added to avoid probabilities equal to 0. Similar to the aff heuristic, this heuristic is also expensive to compute when using the count anonymity measure considered in this paper, as the number of affected nodes need to be computed.
$P_{\text{{ua}}}(\{v,w\})=\frac{|A_{{M}}(\{v,w\})\cap V_{u}|+1/|E|}{\sum_{\{v^{\prime},w^{\prime}\}\in E}(|A_{M}(\{v^{\prime},w^{\prime}\})\cap V_{u}|+1/|E|)}$ (9)
5.4 Utility
Network utility can be defined in many ways. In this work, we choose to measure utility based on the change in topological network properties, and performance on three common downstream tasks in network analysis.
  1. 1.
    Clustering coefficient. The average clustering coefficient over all nodes. This value will likely decrease as triangles are destroyed by deleting edges.
  2. 2.
    Shortest path length. The average shortest path length $d(v,w)$ over all pairs of nodes in the same component. As edges are deleted, paths are destroyed, hence the average shortest path length increases. However, when the network starts to split into components, this value can decrease.
The considered downstream tasks are the following.
  1. 1.
    Network robustness. Change in the fraction of nodes in the largest connected component (LCC). This is expected to decrease as edges are deleted. A more robust network is more resilient under attacks [55].
  2. 2.
    Centrality. Top 100 most central nodes according to betweenness centrality before and after anonymization. If the top 100 nodes did not change the value equals 1.0, and 0.0 if there are no common nodes.
  3. 3.
    Community structure. Measured by the normalized mutual information [51] between the communities [5], before and after anonymization. A value of 1.0 implies the community structure is preserved, and 0.0 that it is completely lost.
6 Results
In this section we first summarize the experimental setup and network datasets used. Second, we continue with two exploratory experiments to compare the effectiveness of different alteration operations on anonymization of graph models, and compare the effect of using different anonymity measures on real-world data. Third, we compare the anonymization algorithms based on their effectiveness on the three variants of the problem. Fourth, we assess the utility achieved by the algorithms on the different variants, and investigate the achieved trade-off between anonymity and utility. Finally, we compare the runtimes of the algorithms on the different variants.
@@ 6 Results
Table 2: Real-world networks used for experiments and their properties: the number of nodes, edges, average degree, clustering coefficient, fraction of nodes in the giant component (LCC), diameter and average distance. Results: initial uniqueness using count as anonymity measure and improvement ratio of ua compared to es. Initial uniqueness values lower than 0.05, i.e., for partial anonymous networks, are shown in italics.
Improvement UA vs. ES
$|V|$ $|E|$
@@
$U(G)$
1.
Full
2.
Partial
3.
Bud-
geted
Radoslaw emails [56] 167 3,250 38.92 0.69 1.00 5 1.97 0.766 8.0 2.3 1.9
Primary school [57] 242 8,317 68.74 0.53 1.00 3 1.73 0.975 4.5 1.8 1.0
Moreno innov. [56] 241 923 7.66 0.31 0.49 5 2.47 0.245 6.2 1.8 2.1
Gene fusion [56] 291 279 1.92 0.00 0.08 9 3.90 0.024 4.8 1.0 5.6
Copnet calls [58] 536 621 2.32 0.25 0.65 22 7.37 0.024 5.5 1.0 8.6
Copnet sms [58] 568 697 2.45 0.22 0.80 20 7.32 0.026 2.4 1.0 3.1
Copnet FB [58] 800 6,418 16.05 0.32 1.00 7 2.98 0.488 7.6 2.1 1.4
FB Reed98 [59] 962 18,812 39.11 0.33 1.00 6 2.46 0.778 5.3 2.3 1.0
Arenas email [56] 1,133 5,451 9.62 0.25 1.00 8 3.61 0.230 8.4 1.7 2.4
Euroroads [56] 1,174 1,417 2.41 0.02 0.03 62 18.37 0.003 1.8 1.0 3.7
Air traffic control [56] 1,226 2,408 3.93 0.07 1.00 17 5.93 0.042 5.4 1.0 5.6
Network science [56] 1,461 2,742 3.75 0.88 0.00 17 5.82 0.039 10.0 1.0 44.4
FB Simmons81 [59] 1,518 32,988 43.46 0.33 0.99 7 2.57 0.785 2.4 2.2 1.4
DNC emails [56] 1,866 4,384 4.70 0.59 0.98 8 3.37 0.092 15.5 1.5 1.4
Moreno health [56] 2,539 10,455 8.24 0.15 1.00 10 4.56 0.054 6.7 1.0 3.6
US power grid [56] 3,783 14,124 7.47 0.28 1.00 10 3.57 0.008 5.2 1.0 3.3
Bitcoin alpha [59] 4,941 6,594 2.67 0.11 1.00 46 18.99 0.113 66.5 1.5 1.9
GRQC collab. [60] 5,241 14,484 5.53 0.69 0.79 17 6.05 0.054 42.2 1.4 3.1
Pajek Erdős [56] 6,927 11,850 3.42 0.40 1.00 4 3.78 0.044 55.7 1.0 4.0
FB GWU54 [59] 12,193 469,528 77.02 0.22 1.00 9 2.83 0.682 - - 1.1
Enron email [60] 36,692 183,831 10.02 0.72 0.92 13 4.03 0.071 - - 2.0
FB wall 2009 [56] 45,813 183,412 8.01 0.15 0.96 18 5.60 0.033 - - 4.6
Brightkite [59] 58,228 214,078 7.35 0.27 0.97 18 4.92 0.048 - - 2.6